Grade 12Physical Chemistry

2g of metallic carbonate on decomposition gave 1.5g of metallic oxide.the equivalent mass of metal is

Yakamallu
8 Years agoGrade 12

1 Answer

deep saini
8 Years ago
Wt of metal carbonate/wt of metallic oxide =(eq.mass of metal + eq.mass of CO3- with 2- charge)/(eq.mass of metal + eq.mass of O with 2- charge)2/1.5=x + 30/x + 8x=58